CARA Telecon Minutes Date: March 26, 2001 Time: 1:00 PM Attendance: BU: Cornell: G. Stacey CMU: J. Peterson Northwestern: CfA: T. Stark South Pole: C. Martin, C. Kaminski Chicago: J. Carlstom, A. Harper, S. Meyer, N. Odalen R. Pernic, R. Landsberg, R. Loewenstein, J. Bausch UCB: W. Holzapfel UCSB: Approval of previous minutes: S. Meyer made a motion, R. Landsberg seconded Agenda Items News from the Pole: AST/RO is observing. Currently doing a map of the SMC CO 2-1. Data is usable but some small problems. After this is completed ~ 2 weeks AST/RO will switch to Wanda. Per C. Martin - Tau is 4-5, weather is poor at the Pole. 350 micron tipper was used to determine the tau. Summit (AASTO 350 um tipper) as reported by J. Peterson - running and working but the Australians have not looked at data. The do not have the time right now to analyze the data. AASTO is working fine per Michael. C. Kaminski reported that they are still fighting on the temps in MAPO, but weather is warm for now. Everything working smoothly. RPS is helping with MAPO temperatures. Helium - ~ 10 days the lost rate is 15-17 liters per day measured on flow. The pressure is down to 2-2.5. R. Pernic suggested that the pressure should be at the point that the transfer is made so winterovers do not have to also mess with that adjustment. Acbar - failure of Azimuth techron amp. The experiment is off the air right now and we are trying to get a spare amp from DASI. Acbar has been shut down ~ 24 hours, but has already taken a lot of data. We are learning a lot about the instrument. Steve Meyer pointed out that since the current limits are well below Techron spec, that likely the failure was due voltage transients due to cut out switch with inductive load. *** J. Peterson to speak with M. Newcomb on what happened before new amp is installed. Amp is rated for 100 amps and experiment requires no more that 20 amps. T. Stark mentioned that there were some spare amps at AST/RO. They run at 20 amps and 40 volts. AST/RO has 6 spares. J. Carlstrom OK'd use of a DASI amp, but need to assess cause of problem and install fix before installation. DASI - working. Bugs reported at last telecon have been cured. Observing on early season tests taking place. Major milestone has been accomplished on last year's results and the papers should be posted at end of April. Question raised by R. Pernic - size of plume on the AASTO. C. Kaminski reported plume is ~ 1.5 feet tall - looks like Swiss cheese, but not hampering operations. There appears to be no back pressure. Smoke keeps coming out. Stack blaster is ineffective anhd not being used. Instrument wish list - J. Peterson has received lots of suggestions. Total of list 50-100k. We may want to leave off small items. We need to determine that the nature of the equipment being requested is truly items that should be purchased by Raytheon/NSF. CARA should send Brian Stone a copy of this list so he is aware of what is being sent in CARA SIP. *** J. Carlstrom to speak with Brian Stone. Tony S wanted it mentioned that Paul Sullivan borrowed a vacuum gauge from AST/RO, which has never been returned. Raytheon/NSF should get their own. T. Stark suggested that this equipment list should include items that Raytheon should purchase for their own use, i.e., in the Cryo Barn the scale should be equipped and our SIP should not be charged for these items. Raytheon could use some more vacuum items to complete the set of spares. There is no equipment for measuring the RFI for the marisat. Carlstrom mentioned that this should be brought up at SPUC. Noted by C. Martin that the TDRS satellite connection is now fixed - problems were mostly on the North American side.